# Gaussian Herding across Pens:::: An Optimal Transport Perspective on Global Gaussian Reduction for 3DGS

### v1

@@ 5 Conclusion
Qiong Zhang is supported by the National Natural Science Foundation of China Grant 12301391.
@@ A.4 Algorithm Convergence
Appendix B More Background and Involved Techniques in Paper
B.1 Gaussian splatting and efficient variants
Due to its advantages in rendering speed and visual fidelity, 3DGS has been adopted in a broad range of applications, including human reconstruction, AI-generated content, autonomous driving, and beyond [34, 35, 36, 37]. Extensions to dynamic 3DGS, editable 3DGS, and surface representation have further broadened its utility [38, 39, 40]. However, the standard 3DGS suffers from significant storage, memory, and rendering costs when applied to large-scale or complex scenes. This limits its deployment on resource-constrained platforms and in broader areas.
To improve storage and rendering efficiency while maintaining visual quality, recent research has focused on two complementary directions: compression and compaction [8]. Compression methods shrink file size by exploiting redundancy in Gaussian attributes (e.g., using vector quantization to cluster similar parameters) [11, 41] or structuring Gaussians into more compact forms (e.g., grids, anchors, and shared codes) [42, 43]. On the other hand, compaction aims to reduce the number of Gaussians required for accurate rendering by optimizing their spatial distribution and pruning redundant primitives [12, 6, 5]. Hybrid pipelines such as LightGaussian [11] and Octree-GS [7] combine compression and compaction together to deliver highly compact and lightweight neural renders.
B.2 Detailed 3DGS pipline
@@
B.3 Involved 3DGS Variants
Although 3DGS features a well-designed pipeline and an efficient differentiable rasterization scheme, its densification process still suffers from several limitations at the detail level. For example, the non-uniform distribution of Gaussians can lead to overlapping primitives in dense regions and under-reconstruction in sparse areas. In addition, the optimization tends to favor large Gaussians that quickly cover the scene to achieve fast convergence. This behavior, however, often results in poor reconstruction of high-frequency details and introduces geometric noise.
Consequently, a growing body of work has focused on refining the densification stage of the original 3DGS framework. In this section, we highlight two such methods—Atom-GS [13] and Mini-Splatting (-D) [12]—which specifically enhance the densification process. While these methods also modify other components such as pruning, we isolate and utilize only the early-stage point cloud generated before 15K training iterations. Applying our compaction method to these two distinct variants further illustrates its generality.
It is important to note that our approach is broadly applicable to any point cloud produced by 3DGS-style pipelines, as long as each Gaussian retains its mean, covariance, and opacity, along with a reasonable optimization strategy for opacity and color.
Atom-GS
To address the tendency of 3DGS to rely on large Gaussians for rapid scene coverage, AtomGS replaces the free-form Gaussian optimization in 3DGS with a structured formulation based on atom Gaussians—uniform, isotropic primitives that better adhere to the underlying scene geometry. In addition, AtomGS introduces geometry-guided optimization to further enhance the spatial fidelity of the reconstructed scene. AtomGS builds upon two key techniques:
  1. 1.
    Atomized Proliferation.
    • Constrains Gaussians in high-detail regions into fixed-size, isotropic Atom Gaussians.
    • Uses a geometric schedule to refine scale and enhance surface coverage over training iterations.
    • Merges Gaussians in flat areas to reduce redundancy.
  2. 2.
    Geometry-Guided Optimization.
    • Introduces Edge-Aware Normal Loss by weighting normal smoothness using an edge map derived from the RGB image.
    • Preserves sharp features while smoothing flat regions, improving alignment with the true 3D geometry.
Mini-Splatting (-D)
The original 3DGS often suffers from non-uniform Gaussian distributions, resulting in excessive overlap in dense regions and insufficient reconstruction in sparse areas. Mini-Splatting addresses this issue by reorganizing the spatial distribution of Gaussians through a two-stage pipeline consisting of densification and simplification. Its goal is to reduce the total number of Gaussians without compromising rendering quality. Mini-Splatting-D is a variant of Mini-Splatting, tailored for quality-oriented rendering. While it follows a similar overall pipeline, the key difference lies in the omission of the simplification stage, allowing the method to focus entirely on maximizing rendering quality. In our implementation, we focus solely on the densification process used in our pipeline.
  • Blur Split: Identifies Gaussians that dominate blurry regions (via screen-space analysis) and splits them to recover fine detail.
  • Depth Reinitialization: Reprojects depth maps derived from midpoints between ray-ellipsoid intersections to generate dense and uniform 3D points for reinitialization.
Beyond evaluating compatibility with different 3DGS densification variants, we compare our method against recent state-of-the-art compaction techniques. Specifically, we consider two representative baselines: LightGaussian [11] and PUP-3DGS [15]. A brief overview of each method is provided below.
LightGaussian
Original 3DGS overpopulates Gaussians during densification, causing storage bloat and rendering inefficiency. LightGaussian prunes redundant Gaussians based on a global significance score (computed via hit count, opacity, and normalized volume), followed by fine-tuning to recover lost details. The steps include:
  1. 1.
    Compute significance for each Gaussian by a score based on the volume of the Gaussian sphere and the number of covered pixels.
  2. 2.
    Prune low-score Gaussians.
  3. 3.
    Fine-tune retained Gaussians with photometric loss.
PUP-3DGS
LightGaussian employs a visibility-aware importance score, which tends to favor pruning small Gaussians while retaining larger background primitives. In contrast, PUP-3DGS introduces a mathematically principled importance measure based on second-order derivatives, allowing it to preserve small but geometrically critical Gaussians in foreground regions. The technical steps are:
  1. 1.
    Fisher Approximation. Computes the gradients of the Gaussian parameters to approximate the Hessian matrix, which is then used to derive the sensitivity score.
  2. 2.
    Iterative Pruning. Rank Gaussians by sensitivity scores, prune low-sensitivity ones, and fine-tune iteratively to recover lost fidelity.
Appendix C Sensitivity Analysis
C.1 Choice of KD-tree Depth
In our algorithm, the KD-tree depth $d$ (or equivalently, the number $s$ of Gaussians per KD block) is treated as a hyperparameter. We investigate how rendering quality varies with different values of $d$ on the Truck scene. The results are shown in Figure 6.
Figure 6: Sensitivity Analysis of $d$. Our GHAP algorithm exhibits stable rendering performance across varying KD-tree depths $d$.
We observe that, within a reasonable range, the KD-tree depth $d$ has minimal impact on rendering quality. In our experiments on the Truck scene, setting $d<6$ results in distance matrices that are too large to fit into GPU memory, causing memory overflow. Conversely, when $d>16$, the number of partitions becomes excessively large, and most blocks contain one or zero Gaussians, rendering the algorithm inapplicable.
C.2 Sampling Iteration
Since most 3DGS-based algorithms complete the densification stage before 15K training iterations, we apply our sampling procedure after this point to achieve a target retention ratio. To study the effect of sampling at different post-densification stages, we evaluate our method at various iterations beyond 15K. The experimental results are presented in Figure 7.
Figure 7: Sensitivity Analysis of Sampling Iteration. The rendering performance of our GHAP algorithm has negligible loss by variations in sampling iteration.
As shown in the results, sampling at any iteration beyond 15K has minimal impact on the final rendering quality. These results indicate that continuing optimization beyond 15K provides diminishing returns. Moreover, once sampling is performed, only a small number of fine-tuning iterations are needed for the compacted model to converge. Based on these observations, we recommend applying sampling uniformly at the 15K iteration, followed by an additional 15K fine-tuning iterations for all experiments.
Appendix D Additional Scene Visualizations
We report a more comprehensive set of results for the plug-in experiments in Table 4, including two additional metrics: the number of Gaussians (in thousands) and rendering speed measured in FPS. As shown in the updated results, our compaction method significantly improves FPS compared to the original models, often yielding over a 2$\times$ speedup in rendering performance.

### v2

@@ 5 Conclusion and Discussion
Qiong Zhang is supported by the National Natural Science Foundation of China Grant 12301391. The authors would like to thank the anonymous reviewers for their constructive suggestions.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Our main claim are summarized in Section 1 and Figure 1, Section 3 and Section 4 offer detailed explainations.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We include the limitations of our work in Appendix.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We have stated all assumptions explicitly within the theorem and include the complete proof of our theoretical results in Appendix.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We explained our settings and hyperparameters in Section 4 and Appendix .
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The data is open-source and the GitHub link is provided.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We placed the detailed experimental settings of the experiments in Section 4 and the Appendix of the paper and the complete details with the code in the supplementary materials.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We show results in Appendix.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: All experiments were conducted on a server with 256 GB RAM and 64 cores Intel ® Xeon ® Gold 6330 CPU, and a standard workstation equipped with a single NVIDIA RTX 3090 GPU.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research conducted in this paper complies with the NeurIPS Code of Ethics in all aspects.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: This paper proposes a novel method of compaction for 3D Gaussian Splatting, which is broadly applicable to 3D real-time radiance field rendering. Its potential impact is further analyzed in Section 5.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This paper involves no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We have ensured that all assets used in this paper are properly cited and their owners credited.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: Our implementation code is available in supplementary materials.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This paper does not involve crowdsourcing nor human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This paper does not involve crowdsourcing nor human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The development of the core method in this research did not involve LLMs as any important, original, or non-standard components.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
@@ A.4 Algorithm Convergence
Appendix B Experiment Steps and Complexity Analysis in More Detail
B.1 Detailed 3DGS Pipline
@@
B.2 Algorithm Complexity
For Algorithm 1, we discuss its computational cost for reducing $n$ Gaussians to $m$ Gaussians. The assignment step requires computing pairwise distances between all $n$ input Gaussians and $m$ cluster centers, resulting in a complexity of $\mathcal{O}(nm)$. The update step involves computing the barycenters. Since the assignments are already known and the update for each cluster is linear in the number of assigned Gaussians, this step has a complexity of $\mathcal{O}(n)$. The overall time complexity of Algorithm 1 is ${\mathcal{O}}(nm)$.
Our GHAP algorithm consists of two components: Geometric Compaction and Appearance Optimization (fine-tuning). In this analysis, we focus only on the time complexity of the geometric compaction stage. The KD-tree is constructed by recursively splitting the dataset along the median of a selected coordinate axis until a maximum depth $d=\lfloor\log_{2}(n/s)\rfloor$ is reached. Let the input size be $n$, and the time complexity of building the tree be $T(n)$. This follows the recurrence: $T(n)=2T(n/2)+\mathcal{O}(n\log_{2}n).$ The recursion terminates after $d$ levels, yielding a total KD-tree construction complexity of $\mathcal{O}(nd\log_{2}(n))$. In the second step of geometric compaction, Algorithm 1 is applied independently within each KD-tree block with $s$ Gaussians reduced to $\rho s$ Gaussians. Based on our analysis in the previous paragraph, the per-block cost is $O(\rho s^{2})$. This along with the fact that there are $2^{d}=n/s$ blocks, the total complexity becomes: $\mathcal{O}\left((n/s)\cdot\rho s^{2}\right)=\mathcal{O}\left(\rho ns\right)={\mathcal{O}}(ms).$ Combining with the KD-tree construction cost, the total time complexity of the geometric compaction step is $\mathcal{O}\left(n(d\log n+mT/{2^{d}})\right).$
B.3 Detailed Experiments Steps
To ensure a fair comparison in our experiments, all methods undergo 30,000 total iterations under consistent training conditions. MCMC uses 30K iterations of its own update process in [24]. For other baselines, we summarize the backbone architecture, compaction methods, and fine-tuning steps in the table below.
Method Backbone Compaction Fine-tune
(0-15k iterations) (15001th iteration) (150001-30K iterations)
3DGS a None d
3DGS+GHAP a Our compaction d
LightGaussian a LightGaussian compaction d
PUP-3DGS a PUP-3DGS compaction d
Trimming the FAT a Trimming the FAT compaction d
MesonGS a MesonGS compaction d
MiniSplatting b MiniSplatting compaction d
MiniSplatting-D+GHAP b b Our compaction d
LocoGS c LocoGS compaction d
  • a
    Vanilla 3DGS densification and pruning in [3].
  • b
    Mini-Splatting densification and pruning in [12].
  • c
    LocoGS 3DGS update in [33].
  • d
    3DGS fine-tuning in [3].
Key implementation details:
  • For compaction methods applicable to pre-trained models (e.g., LightGaussian, PUP-3DGS, Trimming the FAT, MesonGS), we initialize from the same backbone model (trained using vanilla 3DGS) and apply their respective compaction directly—excluding any compression modules.
  • All methods undergo identical fine-tuning (15k iterations) to achieve the target retention ratio.
This standardized protocol ensures that performance differences stem solely from the methods themselves, eliminating variations due to training procedures. All experiments are conducted on a server with 256 GB RAM and a 96-core Intel Xeon Platinum 8255C CPU, and on a workstation equipped with five NVIDIA RTX 3090 GPUs, each with 24 GB of VRAM.
Appendix C An Additional Comparison Experiments
As documented in prior work, 3DGS-MCMC [24] reinterpretes the 3D Gaussian Splatting (3DGS) process through the lens of Markov Chain Monte Carlo (MCMC), thereby naturally exploring a broader parameter space. In particular, it formulates pruning as a state transition within the MCMC framework and incorporates $L_{1}$ regularization, avoiding abrupt hard-threshold deletions and enabling the natural removal of unnecessary Gaussian elements. Given these properties, a further comparative analysis with 3DGS-MCMC is warranted.
To facilitate a fair comparison, we configured 3DGS-MCMC under two settings: one constrained to 300k Gaussians, and another initialized with 3000k Gaussians followed by compaction using our method to reduce the count to 300k. Quantitative results on Tandt & Temples are summarized in the table below.
[…]
3DGS-MCMC-300k 0.813 22.925 0.239
3DGS-MCMC-3000k+GHAP 0.827 22.786 0.209
As evidenced by the results presented in the table, GHAP combined with 3DGS-MCMC yields superior performance compared to the native 3DGS-MCMC approach constrained to 300k Gaussians. This demonstrates the effectiveness of our method in accurately approximating the surface geometry of 3DGS, highlighting the advantage of our compaction strategy.
Appendix D Ablation on Joint Geometry and Appearance Fine-tuning
In our configured compaction process, compaction was performed only once at the 15,001st iteration. We conducted an ablation study on the timing of compaction to examine the impact of varying compaction frequencies on the final outcome. Two alternative compaction strategies were considered: the first involved compacting to 20% at the 15,001st iteration, followed by an additional 50% compaction at the 20,001st iteration; the second strategy applied compaction of 40%, 50%, and 50% at the 15,001st, 20,001st, and 25,001st iterations, respectively. All three strategies ultimately resulted in a final retention rate of 10%. Experimental results, as presented in the table below, indicate that jointly refining geometry and appearance over multiple stages does not lead to evidently improved performance. Therefore, in practical applications, emphasis should be placed on executing appearance optimization for as long as possible, rather than pursuing multi-stage compaction.
[…]
15001 0.1 0.817 23.313 0.242
15001, 20001 0.2,0.5 0.817 23.44 0.247
15001, 20001, 25001 0.4,0.5,0.5 0.811 23.359 0.255
Appendix E Additional Numerical Results and Scene Visualizations
We report a more comprehensive set of results for the comparison experiments in Table LABEL:tab:20, including one different retention ratio: 20%. As shown in the updated results, our compaction method significantly outperforms other methods either post-processing compression methods or end-to-end methods.